how do you solve:
A ladder is resting against a wall. The top of the ladder touches the wall at a height of 9 ft. Find the length of the ladder if the length is 3 ft more than its distance from the wall.
Thank you!

Answer by mananth(16946)

A ladder is resting against a wall. The top of the ladder touches the wall at a height of 9 ft. Find the length of the ladder if the length is 3 ft more than its distance from the wall.
..
let distance from wall be x
length of ladder = x+3
height =9
...
(x+3)^2-x^2=9^2
x^2+6x+9-x^2=81
6x+9=81
6x=81-9
6x=72
/6
x=12 ft distance from wall
length of ladder = 12+3=15 ft
...
m.ananth@hotmail.ca